Sall has a tin can that contain $5.65 in quarters and dimes. there are 34 coins in the can. how many of each type does she have?
Answer by mananth(16949)

Sall has a tin can that contain $5.65 in quarters and dimes. there are 34 coins in the can. how many of each type does she have?
...
quarters = q
dimes = d
...
q+d=34.................1
25q+10d=565 cents
/5
5q+2d=113......2
..
multiply (1) by -5
-5q-5d=-170
add it to (2)
5q-5q-5d+2d=113-170
-3d=-57
/-3
d=19
q=34-d
q=34-19
q=15
..
Check
15*25+10*19
..
quarters=15,dimes =10
...
m.ananth@hotmail.ca